OpenStudy (anonymous):

HELP ME! 1. Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) 2. The legs of an isosceles right triangle are 11 cm long. Find the length of the hypotenuse. Round to the nearest tenth if necessary.

OpenStudy (anonymous):

1. Add the x coordinates of both the points and tell me what do you get?

OpenStudy (anonymous):

which ones are the x coords?

OpenStudy (anonymous):

\[\large \color{green}{Midpoints = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})}\]

OpenStudy (anonymous):

See every point consist of 2 points.. (x, y) In this bracket () the first one is x coordinate and second one is y coordinate that is (x, y) If you are given with two points then, (x1, y1) and (x2, y2) The first ones are x coordinates and the second ones are y coordinates.. Now separately add the x and y coordinates of both the points and divide the result by 2 you will have your answer..

OpenStudy (anonymous):

i just started this stuff i dont know how to do it

OpenStudy (anonymous):

there will come 2 Coordinates in place of 2 points in first line..

OpenStudy (anonymous):

\[ = (\frac{-2 -3}{2}, \frac{1+2}{2}) = (\frac{-5}{2}, \frac{3}{2})\]

OpenStudy (anonymous):

so - 3 - 5

OpenStudy (anonymous):

And for second one: \[Hypotenuse = \sqrt { Base^2 + Height^2}\] \[\large H = \sqrt{ 121 + 121} = 11\sqrt{2}\]